OpenStudy (anonymous):

1. Solve the system of equations. y = 2x² – 3 y = 3x – 1 2. How many real number solutions are there to the equation 0 = –3x² + x – 4? (1 point) 3. Solve the equation by completing the square. If necessary, round to the nearest hundredth. x² – 18x = 19 (1 point) 1; 19 1; 19 3; 6 –3; 1

OpenStudy (anonymous):

Equate both : 2x2 - 3 = 3x -1 solve for x and obtain y by replacing x into your equations....

OpenStudy (anonymous):

for the first part u should solve an second degree equation . for part two just find delta . delta equation can be in three form positive , negative and zero

OpenStudy (anonymous):

each of them shows certain number of answer . 2 answer , 1 repeated answer and none answer

OpenStudy (anonymous):

For 2. 0 = –3x² + x – 4 find values for x and you i guess it would be 2 real for 3. complete square : (x-9)2 -19 + 18 = 0 and continue to obtain correct answer.

OpenStudy (anonymous):

x² – 18x = 19 re-write x^2 -18x -19 = 0 find 2 numbers, where the SUM is -18 and the PRODUCT is -19 (x+1)(x-19) = 0 Two solutions x= -1 x= 19

OpenStudy (anonymous):

thanks so much guys! you think you can you help me with these three? 1. Solve. x² – 81 = 0 (1 point) 0 –9 –9, 9 9 2. Graph the quadratic functions y = -2x2 and y = -2x2 + 4 on a separate piece of paper. Using those graphs, compare and contrast the shape and position of the graphs. 3. Which model is most appropriate for the data shown in the graph below? quadratic linear exponential line

OpenStudy (anonymous):



OpenStudy (anonymous):

1.) x² – 81 = 0 x² = 81 √x² = √81 x = 9

OpenStudy (anonymous):

nope chooch you should have 2 values fox x ...

OpenStudy (anonymous):

@ chooch : i think we should consider both +and - for 9as antoni proofed

OpenStudy (anonymous):

ohhh opps! the answer is c.

OpenStudy (anonymous):